Lightfoot Corp. incurred the following expenditures: $8,000 cost of annual property insurance on the company’s warehouse; $12,000 cost to develop and register a trademark; $30,000 cost to pave the parking lot for the company’s distribution center; $2,000 cost of repairs and maintenance on the company’s lawn service equipment. The total amount of these expenditures that Lightfoot Corp. should be capitalized are?

Multiple Choice

  • $38,000.

  • $44,000.

  • $42,000.

  • $50,000.

Moped, Inc. purchased machinery at a cost of $44,000 on January 1, 2020. The expected useful life is 5 years and the asset is expected to have salvage value of $4,000. Moped depreciates its assets using the double-declining balance method.

What is the firm's depreciation expense for the year ended December 31, 2020?

Multiple Choice

  • $17,600

  • $12,000

  • $8,800

  • $4,000
Answer #1

1. Amount to be capitalized

= 12,000+30,000

= 42,000

Option C

.

Double declining rate = 200/useful life = 200/5 = 40%

= 44,000*40%

= 17,600

Option A